### v1

@@ Abstract
Full Waveform Inversion (FWI) is a widely used method in seismic data processing, capable of estimating models that represent the characteristics of the geological layers of the subsurface. Because it works with a massive amount of data, the execution of this method requires much time and computational resources, being restricted to large-scale computer systems such as supercomputers. Techniques such as FWI adapt well to parallel computing and can be parallelized in shared memory systems using the application programming interface (API) OpenMP. The management of parallel tasks can be performed through loop schedulers contained in OpenMP. The dynamic scheduler stands out for distributing predefined fixed-size chunk sizes to idle processing cores at runtime. It can better adapt to FWI, where data processing can be irregular. However, the relationship between the size of the chunk size and the runtime is unknown. Optimization techniques can employ meta-heuristics to explore the parameter search space, avoiding testing all possible solutions. Here, we propose a strategy to use the Parameter Auto Tuning for Shared Memory Algorithms (PATSMA), with Coupled Simulated Annealing (CSA) as its optimization method, to automatically adjust the chunk size for the dynamic scheduling of wave propagation, one of the most expensive steps in FWI. Since testing each candidate chunk size in the complete FWI is unpractical, our approach consists of running a PATSMA where the objective function is the runtime of the first time iteration of the first seismic shot of the first FWI iteration. The resulting chunk size is then employed in all wave propagations involved in an FWI. We conducted tests to measure the runtime of an FWI using the proposed autotuning, varying the problem size and running on different computational environments, such as supercomputers and cloud computing instances. The results show that applying the proposed autotuning in an FWI reduces its runtime by up to 70.46% compared to standard OpenMP schedulers.
@@ 1 Introduction
Seismic surveying is one of the leading geophysical techniques. This technique is used to identify subsurface features, with the ultimate goal of obtaining an image of the subsurface. This method is widespread in the exploration of oil and gas. Input data used in seismic surveying are obtained through seismic acquisition. In this stage, seismic shots, typically near the surface, generate seismic waves that propagate through the subsurface and can be reflected. Sensors, such as geophones or hydrophones, can record the energy of such waves as seismic traces. The collection of such traces is called observed data. They are processed to obtain information from the subsurface.
Seismic modeling can be one of the processing steps. It is responsible for simulating the propagation of waves through the subsurface. This simulation generates information about how seismic waves propagate through the subsurface and how they are affected by the geophysical properties of the subsurface, such as the velocity of seismic wave propagation. The simulation is based on geophysical models of the subsurface developed based on the observed data and other geological information. Many aspects, such as density, acoustic pressure, and particle velocity, can be modeled.
The more modeled parameters, the more accurate the seismic modeling is. On the other hand, increasing the number of modeled parameters can make the computation more complex and time-consuming. The reason is that the simulation of seismic wave propagation involves intensive mathematical calculations that require a lot of computational processing power. The information generated from seismic modeling is called calculated data.
@@
The procedure to obtain a velocity model from seismic data is called velocity analysis. Full Waveform Inversion (FWI) (Tarantola, 1984) is one of the most used methods to estimate a velocity model. An FWI iteratively adjusts the velocity model using an optimization strategy. It aims to reduce the residual data, i.e., the difference between the observed and calculated data. FWI computes adjustments to the velocity models from the calculated and residual data.
@@
FWI also presents challenges, such as the need for high-quality data and the computational complexity of the process. Due to its high computational demands, the execution of this algorithm in industry models is restricted to large computing environments where the use of parallel programming techniques is necessary.
@@
OpenMP (Open Multi-Processing ) (Dagum and Menon, 1998), a library for parallel programming in shared memory systems, is a widely used tool in the parallelization of large-scale applications such as the FWI. It allows programmers to write parallel code to distribute the work of an application between multiple execution threads. Those threads are then assigned to cores of a shared memory system. By default, in OpenMP, this load is divided equally. However, in complex problems such as FWI, a common challenge is load imbalance, where some parts of the code execute unevenly regular tasks, making some threads complete their work more quickly than others. The faster threads will become idle while waiting for the other threads to finish. Load imbalance can result in underutilization of the available hardware and, consequently, a reduction in the program’s overall performance.
@@
Here, we propose a method to balance the workload between CPUs on a computing node, avoiding resource underutilization. To do this, we employ the Parameter Auto Tuning for Shared Memory Algorithms (PATSMA) (Fernandes et al., 2024) to automatically adjust the chunk size of OpenMP dynamic scheduler. The main contribution of this work is to adapt PATSMA for FWI and evaluate its performance. We compare our proposal with standard OpenMP schedulers, showing that the proposed autotuning reduces the FWI execution time in different computing environments testing for several input sizes.
@@
The rest of this paper is organized as follows: Section 2 presents the characteristics of our target application, FWI, and the autotuning method we use, PATSMA. It also discusses workload distribution through loop scheduling supported by OpenMP. Next, in Section 3, we detail our FWI application and the proposed autotuning approach. We show the proposed method’s results compared with the standard OpenMP schedulers in Section 4. Section 5 presents the related work, and Section 6 concludes this work.
@@ 2 Background
This section briefly describes our target application, the FWI (Section 2.1), and the autotuning method we use, PATSMA (Section 2.2). It also presents the standard OpenMP loop schedulers we compare to our approach (Section 2.3).
@@ 2.1 FWI
2.2 PATSMA
The Parameter Auto Tuning for Shared Memory Algorithms (PATSMA) is a software that aims to adjust a parameter of a shared memory application using an optimization method. The application programmer can define both the parameter and the optimization method. Although PATSMA is suitable for different optimization methods, its software package includes Coupled Simulated Annealing (CSA) (Xavier-de Souza et al., 2010).
Coupled Simulated Annealing (CSA) is a global optimization method that can find optimal solutions for problems with multiple local minima. CSA is an extension of the Simulated Annealing (SA) algorithm (Kirkpatrick et al., 1983) that uses multiple SA instances to explore the solution space simultaneously. In CSA, two or more optimizers are coupled, i.e., at each iteration, the optimizers exchange information about the best solution.
The SA algorithm is summarized in two phases: the generation of a new solution and the acceptance of that solution. Some parameters control both process stages, known as the generation and acceptance temperatures. A generation temperature function obtains new candidate solutions. When a candidate solution has a lower cost than the previous solution, it is accepted with a probability of $100\%$. Otherwise, that solution is accepted as the new solution with a probability calculated using an acceptance temperature function.
CSA increases the search diversity by using multiple SA processes in different subsets of the search space. The exchange of information between the optimizers helps to diversify the search, allowing CSA to find global solutions more efficiently. Additionally, this exchange enables automatic adjustment of the acceptance temperature, making CSA less dependent on the initial value of this parameter.
The optimizers can start from different starting points in the search space. As a result, CSA can explore the solution space more widely than standard SA. At each iteration, each optimizer generates a new solution. If that solution worsens, it can still be selected according to some defined probability. The algorithm becomes less likely to accept a worse solution as the process continues. CSA can be a very effective technique for finding optimal solutions to global optimization problems. However, precisely tuning the CSA parameters is essential to ensure that CSA can generate satisfactory solutions for specific problems. The advantage of CSA over SA is that CSA is less sensitive to the choice of parameters and can produce solutions of better quality even when the parameters are not tuned optimally.
@@ 2.3 OpenMP schedulers
Applications with heterogeneous iterations workload or running in heterogeneous hardware may perform better using the dynamic scheduler. Because such a scheduler distributes the chunks of iterations at runtime, no thread stays idle until all the chunks are distributed. However, the program may lose performance if the block size is not chosen appropriately. If the chunk is too large, it can generate idle time. On the other hand, if the chunk is too small, there may be excessive chunk requests, leading to a significant increase in runtime.
The guided scheduler aims to reduce idleness compared to the static scheduler and management overhead compared to the dynamic scheduler. Because the last chunks are smaller, threads will not idle long. Also, as the first chunks are larger, threads will be busy longer, reducing the number of new chunk requests. The disadvantages of the guided scheduler are that (1) there is an additional overhead to manage the chunk size, and (2) an autotuning method could only control the size of the last block.
@@
Assis et al. (2020) evaluated the performance of the dynamic and static schedulers using a variety of simulations in a reverse time migration (RTM) algorithm (Baysal et al., 1983). The results showed that the dynamic scheduler with chunk size autotuning consistently outperformed the standard static and guided schedulers, with improvements of up to $33\%$ in some cases. Therefore, we used the dynamic scheduler in this work.
3 Autotuning applied to FWI
@@
An algorithm to calculate the solution of the 3D wave equation through finite differences comprises nested loops representing the space domain. The solution of the wave equation is one of the most time-consuming parts of an FWI and is executed several times. Those loops occur in FWI in the direct propagation, reverse propagation, and gradient calculation steps. They are often parallelized in shared memory systems using OpenMP API and employing an OpenMP loop scheduler, as shown in Section 2.3. Aspects such as memory size, number of cores, and input size may influence the choice of an ideal scheduler.
@@
To minimize FWI’s execution time, we propose using PATSMA with CSA to adjust the chunk size for OpenMP dynamic scheduling of parallel loops in an FWI. We use the time to solve the 3D wave equation as PATSMA’s cost function and the chunk size for OpenMP dynamic scheduling as its parameter.
Our FWI implementation is based on Assis et al. (2019). This version of FWI uses optimal checkpointing (Symes, 2010) with the library proposed by Griewank and Walther (2000). As optimal checkpointing also computes the solution of the wave equation, we also use the adjusted chunk size to schedule checkpointing loops dynamically.
@@
Algorithm 1 Full Wavefield Inversion with the proposed autotuning
@@ 4.1 Computing Environments
We used different computational environments for the performance tests of the proposed autotuning, as described below:
  • OPT3 - VM.Optimized3.Flex is an Oracle virtual machine with an Intel Xeon 6354 processor with 18 cores, a base frequency of 3.0 GHz, a maximum turbo frequency of 3.6 GHz, and 256 GB of RAM.
  • SD - represents a node of SDumont’s supercomputer located at the National Center for Scientific Computing (LNCC). This node contains 2 Intel Xeon E5-2695v2 Ivy Bridge CPUs (12c @2.4GHz) totaling 24 cores and 64 GB of RAM.
  • NPAD - NPAD represents a node of the NPAD supercomputer located at the Federal University of Rio Grande do Norte (UFRN). This node contains 2 Intel Xeon CPU E5-2683 v4 @ 2.10GHz CPUs, for a total of 32 cores, and 512 GB of RAM
  • STD3 - VM.Standard3.Flex is an Oracle virtual machine with an Intel Xeon Platinum 8358 processor with 32 cores, a base frequency of 2.6 GHz, a maximum turbo frequency of 3.4 GHz, and 512 GB of RAM.
  • STDE4 - VM.Standard.E4.Flex is an Oracle Cloud virtual machine with an AMD EPYC 7J13 processor with 64 cores, a base frequency of 2.55 GHz, a maximum boosting frequency of 3.5 GHz, and 1024 GB of RAM.
  • DENSE - BM.DenseIO.E4.128 is an Oracle Cloud virtual machine with an AMD EPYC 7J13 processor with 128 cores, a base frequency of 2.55 GHz, a maximum boosting frequency of 3.5 GHz, and 2048 GB of RAM.
@@ 4.2 Test parameters
In terms of the FWI parameterization, for all tests conducted here, we used a peak frequency ($f_{peak}$) of $10Hz$, a time sampling interval of $1ms$, a total of $2458$ time steps, a spatial resolution of $\Delta x1=\Delta x2=\Delta x3=10m$, and an absorbing boundary thickness of $25$ grid points in all directions of the 3D mesh. We constructed $c(x)$ using a sphere with a Gaussian perturbation, where the upper layers have velocities of $2500m/s$ and its core has a maximum velocity of $3500m/s$. We used $3$ velocity models of dimensions $(n1,n2,n3)=(100,400,400)$, $(200,400,400)$, and $(400,400,400)$ where $n1$, $n2$, and $n3$ are the number of samples for the spatial dimensions $x1$, $x2$, and $x3$. The model with $n1=400$ is shown in Fig. 1. The others are slices taking the first $n1$ points of the model from Fig. 1 in the $x1$ dimension.
@@
This paper uses CSA’s parameters discussed and tested by Assis et al. (2020). Table 1 contains the parameters used in our tests.
@@ 4.3 Performance analysis
The first set of tests aims to verify if it is possible to reuse the chunk of the first shot for the next ones. The model size chosen was $(n1,n2,n3)=(200,400,400)$ and we used $1$, $2$, $4$, $8$ and $16$ shots. These tests were performed on the NPAD machine, and each value presented represents a median of $5$ points.
Fig. 2 shows that, in all cases, the proposed strategy presented better performance, with speedups of up to $7.98\%$. Increasing the number of shots, we observe that the proposed autotuning performance remains stable because its overhead does not increase, being similar to the first shot. In this set of tests, the overhead of the proposed method remained below $1.2\%$ in all cases. The stable speedup indicates that the chunk size adjusted for the first shot can be reused in the following shots.
@@
Figure 2: Speedup for FWI using the proposed autotuning compared to the OpenMP static and guided schedulers on the NPAD machine, with $1$, $2$, $4$, $8$, and $16$ shots. For the OpenMP schedulers, the chunk size was not explicitly set. The velocity model size for these tests was $(n1,n2,n3)=(200,400,400)$. Each point is a median of five executions.
In another set of tests, we measured the performance of the proposed strategy as we increased the problem size and computational resources. Fig. 3 illustrates the performance of the proposed autotuning compared to the standard OpenMP schedulers static and guided on the set of six computational environments mentioned in Subsection 4.1, varying $n1=\{100,200,400\}$. For the standard OpenMP schedulers, the chunk size was not explicitly set. Each point in the graph is a median of five samples. Fig. 3 shows that the proposed method had a superior performance in all scenarios. It also shows that, in most cases, the more cores a machine has, the higher the obtained speedup. The reason is that the higher the number of cores, the smaller the search space, implying fewer possible solutions (see Section 3). That way, CSA is more likely to find a near-optimal solution.
@@
Figure 3: Single-shot FWI speedup using the proposed autotuning compared to OpenMP (a) static and (b) guided schedulers on $5$ machines (OPT3, SD, NPAD, STD3, STDE4, and DENSE), for three input sizes, $(n1,n2,n3)=(100,400,400)$, $(200,400,400)$, and $(400,400,400)$. For OpenMP schedulers, the chunk size was not explicitly set. Each point is a median of at least five runs.
@@
As we can see, the speedups related to the static scheduler (Fig. 3a) are higher than those related to the guided scheduler (Fig. 3a). A poor performance of the static scheduler for large problem sizes can explain that phenomenon. The larger the problem size for the default static scheduler, the larger its chunk size (see Section 2.3). That means reduced data locality and, consequently, reduced cache reuse. Compared to the static scheduler, the FWI using the proposed strategy obtained a speedup of up to $70.46\%$.
@@
On the other hand, guided scheduler reduces threads’ idle time. Chunks are distributed dynamically, and by the end of the execution, chunk sizes are smaller. That way, the maximum idle time a thread may have is the time for processing the last chunk. guided scheduler can be more efficient than dynamic scheduler in cases where the ideal block size is unknown. However, the wave propagation effect can affect its performance. Should the first chunk (the largest) contain most of the non-zero wavefield, it will dictate the performance. In addition, guided scheduler has an overhead in managing the block size, which does not occur with dynamic scheduler. FWI using the proposed approach had a speedup of up to $15.27$% compared to using guided.
In our last set of experiments, we compared the runtime of an FWI using the proposed autotuning, the OpenMP static and guided scheduler with their default chunk size values, and the OpenMP dynamic scheduler with the chunk size defined by Mohammed et al. (2022):

### v2

@@ Abstract
Full Waveform Inversion (FWI) is a widely used method in seismic data processing, capable of estimating models that represent the characteristics of the geological layers of the subsurface. Because it works with a massive amount of data, the execution of this method requires much time and computational resources, being restricted to large-scale computer systems such as supercomputers. Techniques such as FWI adapt well to parallel computing and can be parallelized in shared memory systems using the application programming interface (API) OpenMP. The management of parallel tasks can be performed through loop schedulers contained in OpenMP. The dynamic scheduler stands out for distributing predefined fixed-size chunk sizes to idle processing cores at runtime. It can better adapt to FWI, where data processing can be irregular. However, the relationship between the size of the chunk size and the runtime is unknown. Optimization techniques can employ meta-heuristics to explore the parameter search space, avoiding testing all possible solutions. Here, we propose a strategy to use the Parameter Auto-Tuning for Shared Memory Algorithms (PATSMA), with Coupled Simulated Annealing (CSA) as its optimization method, to automatically adjust the chunk size for the dynamic scheduling of wave propagation, one of the most expensive steps in FWI. Since testing each candidate chunk size in the complete FWI is unpractical, our approach consists of running a PATSMA where the objective function is the runtime of the first time iteration of the first seismic shot of the first FWI iteration. The resulting chunk size is then employed in all wave propagations involved in an FWI. We conducted tests to measure the runtime of an FWI using the proposed auto-tuning, varying the problem size and running on different computational environments, such as supercomputers and cloud computing instances. The results show that applying the proposed auto-tuning in an FWI reduces its runtime by up to 70.46% compared to standard OpenMP schedulers.
keywords:
[…]
@@ 1 Introduction
Seismic surveying is a technique that involves utilizing seismic shots to emit waves and sensors to capture data from returning waves. This data is then processed to gather insights about what is in the subsurface. Seismic modeling is a part of the surveying process that can be utilized to simulate wave propagation and subsurface properties. While it can model various parameters for increased accuracy, doing so can result in more complex mathematical calculations that require intensive computational processing and time. The result of this process is what is known as calculated data.
@@
Full Waveform Inversion (FWI) (Tarantola, 1984) is one of the most used methods to estimate a velocity model. An FWI iteratively adjusts the velocity model using an optimization strategy. It aims to reduce the residual data, i.e., the difference between the observed and calculated data. FWI computes adjustments to the velocity models from the calculated and residual data.
@@
FWI presents challenges, such as the need for high-quality data and the computational complexity of the process. Due to its high computational demands, the execution of this algorithm in industry is restricted to large computing environments where the use of parallel programming techniques is necessary.
@@
OpenMP (Open Multi-Processing ) (Dagum and Menon, 1998), a library for parallel programming in shared memory systems, is a widely used tool in the parallelization of large-scale applications such as the FWI. It allows programmers to write parallel code to distribute the work of an application between multiple execution threads. Those threads are then assigned to cores of a shared memory system. By default, in OpenMP, this load is divided equally. However, in complex problems such as FWI, load imbalance may arise because some parts of the code execute uneven tasks, making some threads complete their work more quickly than others. The faster threads will become idle while waiting for the other threads to finish. Load imbalance can result in underutilization of the available hardware and, consequently, a reduction in the program’s overall performance.
@@
Here, we propose a method to balance the workload among CPUs on a computing node, avoiding resource underutilization. To do this, we employ the Parameter Auto-Tuning for Shared Memory Algorithms (PATSMA) (Fernandes et al., 2024) to automatically adjust the chunk size of the OpenMP dynamic scheduler. The main contribution of this work is to adapt PATSMA for FWI and evaluate its performance. We compare our proposal with standard OpenMP schedulers, showing that the proposed auto-tuning reduces the FWI execution time in different computing environments and tests for several input sizes.
@@
The rest of this paper is organized as follows: Section 2 presents the characteristics of our target application, FWI, and the auto-tuning method we use, PATSMA. It also discusses workload distribution through loop scheduling supported by OpenMP. Next, in Section 3, we detail our FWI application and the proposed auto-tuning approach. We show the proposed method’s results compared with the standard OpenMP schedulers in Section 4. Section 5 presents the related work, and Section 6 concludes this work.
@@ 2 Background
This section briefly describes our target application, the FWI (Section 2.1), shows the standard OpenMP loop schedulers we compare to our approach (Section 2.2), and presents the auto-tuning method we use, PATSMA (Section 2.3).
@@ 2.2 OpenMP schedulers
Applications with heterogeneous iterations workload or running in heterogeneous hardware may perform better using the dynamic scheduler. Because such a scheduler distributes the chunks of iterations at runtime, no thread stays idle until all the chunks are distributed. However, the program may lose performance if the chunk size is not chosen appropriately. If the chunk is too large, it can generate idle time. On the other hand, if the chunk is too small, there may be excessive chunk requests, leading to a significant increase in runtime.
The guided scheduler aims to reduce idleness compared to the static scheduler and management overhead compared to the dynamic scheduler. Because the last chunks are smaller, threads will not idle long. Also, as the first chunks are larger, threads will be busy longer, reducing the number of new chunk requests. The disadvantages of the guided scheduler are that (1) there is an additional overhead to manage the chunk size, and (2) an auto-tuning method could only control the size of the last block.
@@
Assis et al. (2020) evaluated the performance of the dynamic and static schedulers using a variety of simulations in a reverse time migration (RTM) algorithm (Baysal et al., 1983). The results showed that the dynamic scheduler with chunk size auto-tuning consistently outperformed the standard static and guided schedulers, with improvements of up to $33\%$ in some cases. Therefore, we used the dynamic scheduler in this work.
2.3 PATSMA
The Parameter Auto-Tuning for Shared Memory Algorithms (PATSMA) is a software that aims to adjust a parameter of a shared memory application using an optimization method. The application programmer can define both the parameter and the optimization method. Although PATSMA is suitable for different optimization methods, we used Coupled Simulated Annealing (CSA) (Xavier-de Souza et al., 2010), one of PATSMA’s built-in optimizers.
CSA is an extension of the Simulated Annealing (SA) (Kirkpatrick et al., 1983) algorithm that uses multiple SA instances to explore the solution space simultaneously. CSA increases the search diversity by using multiple SA processes (optimizers) in different subsets of the search space. The exchange of information between the optimizers helps to diversify the search, allowing CSA to find global solutions more efficiently.
The CSA algorithm follows the same steps of generation and acceptance as SA. Both algorithms use temperatures to control their process. However, CSA has an automatic process that controls and diversifies the search by adjusting the acceptance temperature of each optimizer. This allows some optimizers to focus on local search while others on global search.
CSA explores the solution space widely with different starting points. It generates a new solution at each iteration, which can still be accepted even if it worsens. It is effective for global optimization, but proper parameter tuning is crucial. CSA is less sensitive than SA to parameter choice and produces better quality solutions even with suboptimal tuning.
PATSMA auto-tuning is a process that involves iterative loops where a parameter affects the runtime of a specific code section within the loop. To perform the auto-tuning, the target application needs to define the code section where PATSMA will act. However, sometimes the loop progress can negatively influence the auto-tuning as the section’s behavior can change during the progress. To mitigate this issue, a copy of this section can be created before the loop where we will execute the auto-tuning.
Once the section is defined, PATSMA uses the CSA to generate probable solutions for this section’s use and measures the resulting runtime. After all PATSMA iterations are complete, the final solution can be sent to the original section code performer for execution.
3 Auto-tuning applied to FWI
@@
Performing Full Waveform Inversion (FWI) can employ an algorithm to solve the 3D wave equation through finite differences. The algorithm involves nested loops that represent the space domain, and it is executed repeatedly during FWI. These loops are in the direct propagation, reverse propagation, and gradient calculation steps. They are often parallelized using the OpenMP API and an OpenMP loop scheduler. In Section 2.2, we show the need to define the chunk size parameter to improve the loop scheduler’s performance. In some cases, the choice of a chunk may depend on factors such as memory size, number of cores, and input size.
@@
The Full Waveform Inversion (FWI) is an intricate algorithm that involves a varying amount of calculations depending on the space region where the calculations are made. This behavior affects the chunk size definition, which can influence FWI’s execution time. To minimize the time to run an FWI, we propose using PATSMA with CSA to adjust the chunk size for OpenMP dynamic scheduling of parallel loops in an FWI. We use the runtime to solve the 3D wave equation as PATSMA’s cost function and the chunk size for OpenMP dynamic scheduling as its parameter.
Our implementation of FWI is based on work from Fernandes et al. (2025). This version of FWI employs optimal checkpointing (Symes, 2007) using the library proposed by Griewank and Walther (2000). Optimal checkpointing operates by storing only a subset of the intermediate simulation states during the forward propagation stage. During the reverse stage, the states that were not stored are recomputed from the available checkpoints. This method significantly reduces memory requirements, with a logarithmic increase in computational cost relative to the total number of steps. Additionally, since optimal checkpointing also computes the solution of the wave equation, we dynamically adjust the chunk size to efficiently schedule checkpointing loops.
@@
Algorithm 1 Full Wavefield Inversion with the proposed auto-tuning
@@ 4.1 Computing Environments
We used different computational environments for the performance tests of the proposed auto-tuning, as described below:
  • OPT3 - VM.Optimized3.Flex is an Oracle virtual machine with an Intel Xeon 6354 processor with 18 cores, a base frequency of 3.0 GHz, a maximum turbo frequency of 3.6 GHz, 256 GB of RAM, and 39 MB of L3 cache.
  • SD - represents a node of SDumont’s supercomputer located at the National Center for Scientific Computing (LNCC). This node contains 2 Intel Xeon E5-2695v2 Ivy Bridge CPUs (12c @2.4GHz) totaling 24 cores, 64 GB of RAM, and 30 MB of L3 cache.
  • NPAD - NPAD represents a node of the NPAD supercomputer located at the Federal University of Rio Grande do Norte (UFRN). This node contains 2 Intel Xeon CPU E5-2683 v4 @ 2.10GHz CPUs, for a total of 32 cores, 512 GB of RAM, and 40 MB of L3 cache.
  • STD3 - VM.Standard3.Flex is an Oracle virtual machine with an Intel Xeon Platinum 8358 processor with 32 cores, a base frequency of 2.6 GHz, a maximum turbo frequency of 3.4 GHz, 512 GB of RAM, and 48 MB of L3 cache.
  • STDE4 - VM.Standard.E4.Flex is an Oracle Cloud virtual machine with an AMD EPYC 7J13 processor with 64 cores, a base frequency of 2.55 GHz, a maximum boosting frequency of 3.5 GHz, 1024 GB of RAM, and 256 MB of L3 cache.
  • DENSE - BM.DenseIO.E4.128 is an Oracle Cloud virtual machine with an AMD EPYC 7J13 processor with 128 cores, a base frequency of 2.55 GHz, a maximum boosting frequency of 3.5 GHz, 2048 GB of RAM, and 256 MB of L3 cache.
@@ 4.2 Test parameters
In terms of the FWI parameterization, for all tests conducted here, we used a peak frequency ($f_{peak}$) of $10Hz$, a time sampling interval of $1ms$, a total of $2458$ time steps, a spatial resolution of $\Delta x1=\Delta x2=\Delta x3=10m$, and an absorbing boundary thickness of $25$ grid points in all directions of the 3D mesh.
[…]
@@
Table 1 presents the parameters used in our experiments. The selection of the number of CSA optimizers, $m$, is extensively analyzed in Gonçalves-e Silva et al. (2018), where it is shown that the optimal values typically range between $m=4$ and $m=10$ to minimize various functions. In this work, we set $m=4$ to reduce the computational overhead of PATSMA while maintaining the CSA performance. For the initial acceptance temperature, we adopted one of the values suggested in Xavier-de Souza et al. (2010), $T_{0}^{ac}=0.9$. Furthermore, Assis et al. (2020) reports experiments that explored different values for the number of iterations, $N$, and the initial generation, ${T_{0}^{ac}}$. Based on their findings, we selected $N=40$ and ${T_{0}^{ac}}=100$, as these parameters yielded the shortest execution times in all test cases.
@@ 4.3 Performance analysis
The first set of tests aims to verify whether it is possible to reuse the chunk size adjusted for the first shot for the next ones. The chosen model size was $(n1,n2,n3)=(200,400,400)$. Initially, we measured the time per shot for executions with $16$ shots. The tests were conducted on the NPAD machine.
Fig. 2 shows that our proposal achieves better execution times compared to the standard Static and Guided schedulers. The first shot presents an initial overhead due to the computation of the chunk size; however, the performance stabilizes in subsequent executions. The static scheduler exhibits greater variability in execution time across different shots, as it is more sensitive to changes in the computational distribution. In contrast, the Guided scheduler also experiences variations but benefits from a degree of adaptation, mitigating extreme fluctuations. On the other hand, the dynamic scheduler with PATSMA demonstrates higher adaptability, maintaining more stable execution times despite changes in shot distribution.
@@
Figure 2: Runtime per shot for FWI using the proposed auto-tuning compared to the OpenMP static and guided schedulers on the NPAD machine, with $16$ shots. For the OpenMP schedulers, the chunk size was not explicitly set. The velocity model size for these tests was $(n1,n2,n3)=(200,400,400)$.
In another set of experiments aimed at verifying the gain in full execution time per shot, we examined whether the chunk size adjusted for the first shot could be reused in subsequent executions. The chosen model size was $(n1,n2,n3)=(200,400,400)$, and we used $1$, $2$, $4$, $8$, $16$, $32$, $64$, and $128$ shots. These tests were performed on the NPAD machine, with each value representing a median of $5$ points.
Fig. 3 shows that in all cases, the proposed strategy demonstrated better performance, with speedups of up to $14.1\%$ and $14.96\%$ when compared to the standard OpenMP static and guided schedulers, respectively. As we increased the number of shots, we observed that the performance of the auto-tuning strategy remained stable, as its overhead did not increase and remained similar to that of the first shot. In this set of tests, the overhead of the proposed method remained below $1.2\%$ in all cases. We conducted tests with different numbers of shots and observed that PATSMA consistently outperformed the Static and Guided schedulers. As shown in Fig. 2, the execution time for the static scheduler varied significantly depending on the shot, probably due to the influence of the shot positions in the 3D model grid, which affects computational distribution. The Guided scheduler exhibited some level of adaptation, reducing extreme fluctuations but still showing variations. In contrast, PATSMA maintained more stable performance in all cases tested, strengthening its effectiveness in handling variations in shot distribution.
@@
Figure 3: Speedup for FWI using the proposed auto-tuning compared to the OpenMP static and guided schedulers on the NPAD machine, with $1$, $2$, $4$, $8$, $16$, $32$, $64$ and $128$ shots. For the OpenMP schedulers, the chunk size was not explicitly set. The velocity model size for these tests was $(n1,n2,n3)=(200,400,400)$. Each point is a median of five executions.
In another set of experiments, we measured the performance of the proposed strategy as we increased the problem size and computational resources (number of cores and amount of L3 cache). Fig. 4 illustrates the performance of the proposed auto-tuning compared to the standard OpenMP schedulers static and guided on the set of six computational environments mentioned in Subsection 4.1, varying $n1=\{100,200,400\}$. For the standard OpenMP schedulers, the chunk size was not explicitly set. Each point in the graph is a median of five samples. Fig. 4 shows that the proposed method had a superior performance in all scenarios. It also shows that, in most cases, the more cores a machine has, the higher the obtained speedup. The reason is that the higher the number of cores, the smaller the search space, implying fewer possible solutions (see Section 3). That way, CSA is more likely to find a near-optimal solution.
Figure 4: Single-shot FWI speedup using the proposed auto-tuning compared to OpenMP (a) static and (b) guided schedulers on $5$ machines (OPT3, SD, NPAD, STD3, STDE4, and DENSE), for three input sizes, $(n1,n2,n3)=(100,400,400)$, $(200,400,400)$, and $(400,400,400)$. For OpenMP schedulers, the chunk size was not explicitly set. Each point is a median of at least five runs.
@@
As we can see, the speedups related to the static scheduler (Fig. 4a) are higher than those related to the guided scheduler (Fig. 4b). A poor performance of the static scheduler for large problem sizes can explain that phenomenon. The larger the problem size for the default static scheduler, the larger its chunk size (see Section 2.2). That means reduced data locality and, consequently, reduced cache reuse. Compared to the static scheduler, the FWI using the proposed strategy obtained a speedup of up to $70.46\%$.
@@
On the other hand, guided scheduler reduces threads’ idle time. Chunks are distributed dynamically, and by the end of the execution, chunk sizes are smaller. That way, the maximum idle time a thread may have is the time for processing the last chunk. guided scheduler can be more efficient than dynamic scheduler in cases where the ideal chunk size is unknown. However, the wave propagation behavior may affect its performance. Should the first chunk (the largest) contain most of the non-zero wavefield, it will dictate the performance. In addition, guided scheduler has an overhead in managing the chunk size, which does not occur with dynamic scheduler. FWI using the proposed approach had a speedup of up to $15.27$% compared to using guided.
In our last set of experiments, we compared the runtime of an FWI using the proposed auto-tuning, the OpenMP static and guided scheduler with their default chunk size values, and the OpenMP dynamic, static and guided schedulers with the chunk size defined by Mohammed et al. (2022):